An inductor, L, of 2 H and a resistor, R, of 1000Ω are in series across an a.c. voltage of 10 V, frequency 60 Hz. Calculate the current in r.m.s. flowing,
1
Expert's answer
2020-07-20T14:56:32-0400

According to the Ohm's law, the current amplitude is:


"I = \dfrac{U}{Z}"

where "U = 10V" is the volage amplitude and "Z" is the total impedance of the circuit.

The impedance of the series resistor and inductor is:


"Z = \sqrt{R^2 + (2\pi fL)^2}"

where "f = 60Hz" is the frequency of a.c. voltage.

Substituting the numerical values, get:


"I = \dfrac{U}{\sqrt{R^2 + (2\pi fL)^2}} = \dfrac{10}{\sqrt{1000^2 + (2\pi\cdot 60\cdot 2)^2}} \approx 0.00798A"

The r.m.s. of a sinusoidal current is:


"I_{rms} = \dfrac{I}{\sqrt{2}}\approx 0.00564A"

Answer. 0.00564 A = 5.64 mA.